OpenStudy (anonymous):

The Cool Company determines that the supply function for its basic air conditioning unit is S(p) = 30 + 0.006p3 and that its demand function is D(p) = 150 - 0.12p2, where p is the price. Determine the price for which the supply equals the demand.

OpenStudy (anonymous):

like this? \[30 + 0.006p^3=150 - 0.12p^2\]

OpenStudy (anonymous):

@satellite73 Yeah i tried it but when i factored i got p=119980

OpenStudy (anonymous):

@satellite73 i think the number is too big..and it not the onthe the options

OpenStudy (anonymous):

yeah it is way too big

OpenStudy (anonymous):

i have no idea how to solve it because it is a cubic equation i think you need some technology http://www.wolframalpha.com/input/?i=30+%2B+0.006p^3%3D150+-+0.12p^2

OpenStudy (anonymous):

you could start with \[30000+6p^3=150,000-120p^2\]

OpenStudy (anonymous):

it doesn't factor easily though, so i am not sure what you tried